# Talking Heads: Understanding Inter-layer Communication in Transformer Language Models

### v1

@@ 5.2 Intervention Experiments
We traverse the 3D space spanned by the top three inhibition head components (7.9.6, 8.6.2, and 8.6.10) and measure how this changes where the mover head attends, and what the model’s final prediction is [^3] We test in increments of 10 from [-100, 100] along each axis, including every combination. it’s possible this is not the optimal range. In Figure 5 (middle), we visualize this traversal with one dot representing a point in the space that we query. We run the entire dataset with the inhibition components set at this point, and the color represents the index of the object that the mover head attends most to. We find that structure emerges as we add items, where areas of this space represent the first and last object, and wedges fill in the space for each item that gets added. Eventually (around 9 or 10 items) these wedges get small and start to fracture (Figure 5, middle bottom). We connect these two phenomena to the performance of the model: the bias to predict later objects, and the inability to handle longer lists. We believe the model is not capable of traversing this space well enough on its own, even though it learned to represent it, and longer lists cause worse performance because the space gets fragmented into smaller and smaller areas that repersent each item.
@@ 6 Related Work
The SVD has been used for interpretability and weight based analysis in the past [Sule et al., 2023, Praggastis et al., 2022, Martin et al., 2021, beren and Black, 2022].
@@ 7 Discussion & Conclusion
Appendix A The Composition Score with and without Weight Decomposition
@@
Appendix B Limitations
Appendix C Duplicate Token Heads
@@
Appendix D More IOI Interventions
@@
Additional interventions testing the efficacy of inhibition heads to change behavior of a downstream mover head are provided in Figures 8 and 9. We include the remaining inhibition heads and additional control heads for the component scaling experiment from Section 4 in Figure 10
@@
Figure 10: We intervene on the forward pass of the model by replacing the output of some attention head as the vector obtained by scaling a component vector by some scalar $\alpha$. By doing so, the actual attention head pattern has no effect on the downstream performance. We show the inhibition component vectors have the unique effect of controlling the position of the name being attended to by the downstream mover head (9.9). Random control head components do not have this effect.
Appendix E Laundry List Data Generation
@@
Appendix F Inhibition Mechanism in Pythia and Training Progression of Inhibition
@@ F.1 Path Patching on IOI
F.2 Training Progression of Inhibition Components
@@
Appendix G Static Weight Analysis for Circuit Discovery
@@
Appendix H More Information on Composition
H.1 Value Composition: Duplicate Token Heads
@@
H.2 Query Composition: Inhibition Heads
@@
H.3 Key Composition: Induction Heads
@@
Appendix I Extra Laundry List Interventions
The results for scaling individual components across Laundry List datasets (varying number of objects) are in Figures 16, 17, 18, 19, 20, 21, 22, 23, and 24

### v2

@@ 5.2 Intervention Experiments
We traverse the 3D space spanned by the top three inhibition head components (7.9.6, 8.6.2, and 8.6.10) and measure how this changes where the mover head attends, and what the model’s final prediction is [^4] We test in increments of 10 from [-100, 100] along each axis, including every combination. it’s possible this is not the optimal range. We leave out 7.3 because we found that it changes inhibition performance the least (Figure 2), and visualizing with only 3 dimensions is much simpler. In Figure 5 (middle), we visualize this traversal with one dot representing a point in the space that we query. We run the entire dataset with the inhibition components set at this point, and the color represents the index of the object that the mover head attends most to. We find that structure emerges as we add items, where areas of this space represent the first and last object, and wedges fill in the space for each item that gets added. Eventually (around 9 or 10 items) these wedges get small and start to fracture (Figure 5, middle bottom). We connect these two phenomena to the performance of the model: the bias to predict later objects, and the inability to handle longer lists. We believe the model is not capable of traversing this space well enough on its own, even though it learned to represent it, and longer lists cause worse performance because the space gets fragmented into smaller and smaller areas that repersent each item.
@@ 6 Related Work
The SVD has been used for interpretability and weight based analysis in the past [Sule et al., 2023, Praggastis et al., 2022, beren and Black, 2022]. Martin et al. [2021], for example, use the SVD on weight matrices to measure weiht quality to predict generalization performance and whether or not a model is well-trained. The SVD has been used in the training of LoRA modules [Hu et al., 2022], as well as in other finetuning methods [Bałazy et al., 2024, Feng et al., 2024, Wang et al., 2024, Sun et al., 2024, Guo et al., 2024, Karimi Mahabadi et al., 2021]. Our work analyzes model properties that may facilitate these methods to work. Aghajanyan et al. [2021] is perhaps the first to report that pretrained models have a very low intrinsic dimensionality, which helps explain and support our claim that we see so many low rank communication channels in large models. We are therefore excited about the connections between this line of work and work on the linear representations in language models [Elhage et al., 2022, Park et al., 2023], which argues that features in LMs are represented as directions in space. There has been recent work in studying interactions between components in finer-grained feature spaces in the past [Kim et al., 2018, Marks et al., 2024, Geiger et al., 2023, Lepori et al., 2023, Zhang et al., 2024], whereas our approach begins by first analyzing the weights to find substantial connections in subspaces without requiring any data. Future work could connect these two, to find circuits, subnetworks, and/or directions in space for certain behaviors informed by connections in the weights, which alleviates the concern of finding these things from scratch. We outline preliminary results for such an approach for novel circuit discovery in Appendix I.
@@ 7 Discussion & Conclusion
Appendix A Limitations
Appendix B Larger Models’ Performance When Increasing the # of Objects in the Laundry List Task
In the main text, we consider models that are quite small by today’s standards, and it is reasonable to wonder whether the inability to handle more objects in context goes away with model scale. In Figure 6, we show that even up to 6.9B parameters, OPT [Zhang et al., 2022] and Pythia models also struggle with increasingly many objects, although degrade more slowly with scale. Regardless, we are not claiming that we have identified a problem that can not be solved through scaling, nor is that of particular interest in this work. Increased capacity is likely to yield mechanisms with higher capacity to handle more objects. What is of greater interest for future work would be to identify whether larger models use fundamentally different mechanisms than the one we identified here to solve the task.
Figure 6: Larger models also degrade performance rapidly as we increase the number of objects in the Laundry List task, although more slowly than smaller models. Pythia 6.9B retains strong performance up to around 10 objects.
Appendix C The Composition Score with and without Weight Decomposition
@@
Appendix D Duplicate Token Heads
@@
Appendix E More IOI Interventions
@@
Additional interventions testing the efficacy of inhibition heads to change behavior of a downstream mover head are provided in Figures 9 and 10. We include the remaining inhibition heads and additional control heads for the component scaling experiment from Section 4 in Figure 11
@@
Figure 11: We intervene on the forward pass of the model by replacing the output of some attention head as the vector obtained by scaling a component vector by some scalar $\alpha$. By doing so, the actual attention head pattern has no effect on the downstream performance. We show the inhibition component vectors have the unique effect of controlling the position of the name being attended to by the downstream mover head (9.9). Random control and head components with other functions (like 8.3.0) do not have this effect.
Appendix F Laundry List Data Generation
@@
Appendix G Inhibition Heads Behavior on Open-domain Text
We find that inhibition heads are consistently active on tokens about to predict the continuation of a sequence (“serotonin, dopamine, and…") and attend to previous items in that sequence, consistent with their role in IOI and Laundry List. We therefore argue that the role the inhibition mechanism plays in both IOI and the Laundry List task performs this same operation in generic language modeling
Figure 12: Examples of high attention in examples in OpenWebText-10K for head 8.6 in GPT-2
Figure 13: Examples of high attention in examples in OpenWebText-10K for head 8.10 in GPT-2
Figure 14: Examples of high attention in examples in OpenWebText-10K for head 7.9 in GPT-2
Figure 15: Examples of high attention in examples in OpenWebText-10K for head 7.3 in GPT-2
Appendix H Inhibition Mechanism in Pythia and Training Progression of Inhibition
@@ H.1 Path Patching on IOI
H.2 Training Progression of Inhibition Components
@@
Appendix I Circuit Discovery with Static Weight Analysis
One of the core claims of this work is that we can find meaningful connections between attention heads by reading them off of the weights. In this section we extend this idea beyond the IOI circuit and use our method to show that we can find novel meaningful connections without running the model. First, we show that we can rediscover the IOI circuit directly from the weights using the decomposed composition score (without knowing anything about its function). Then, we show that we can identify new connections to the 9.9 mover head that is part of a different circuit. We finally run the model on open-domain data and perform causal experiments to begin to identify the functionality of this new circuit, which we implicate in retrieving relevant information from context. While the experiments we perform to identify functionality are not a replacement for a full circuit analysis, we at least show how our method can be used to pinpoint circuits in a network for further analysis. Future work can extend the analysis of the context-retrieval circuit we find here.
I.1 Finding the IOI Circuit in GPT-2 Weights
@@
I.2 Discovering New Connections
We now take this further to show that we can use the composition score to facilitate the discovery of new circuits directly from the weights. Although we can not yet identify function directly from the weights, we can identify components that communicate strongly (indicated by their composition scores) to know where to look for circuits that may explain other behaviors in the model.
We focus on one specific extension to the inhibition-mover subcircuit by looking for new connections into the 9.9 mover head QK circuit. We search for connections into this head because it seems likely that other heads besides inhibition heads control what the mover head attends to, but it has not been investigated in prior work. We look for component matrices that have a high composition score with the 9.9 QK matrix, and plot a heatmap of the highest composing component matrix per head in Figure 23. We find that there are several heads that have components that highly compose with 9.9 that have not been previously identified as affecting copying behavior. These are heads 6.1.1, 6.2.5, 6.7.3, and 8.3.0. We show the per component composition scores for these heads in Figure 24 and find that they also contain individual components with very high relative composition compared to other components (with 6.2 components being distributed slightly more smoothly than the others).
I.2.1 Attention Pattern Analysis of 6.1, 6.2, 6.7, and 8.3
We perform a similar attention pattern analysis as shown in Section G on these heads. We run the model on documents from OpenWebText-10K and look at contexts in which these heads attend strongly (.$\geq.5$ of probability mass) to some token. We find that 6.2 is not interpretable in this way, as its attention patterns are more diffuse, but examples for the other heads are shown in Figures 21, 22, and 20. We find qualitatively similar patterns in these heads to inhibition heads, though with less selection for attending to previous items in lists. Head 8.3 is of particular interest because it has the strongest composition with 9.9, and we will focus on it for the remainder of this section. This head also has some particularly salient motifs in its attention patterns: much of the time it appears to attend to a token (or the last token in a multitoken phrase) in context that could be a plausible continuation of the current text. In the following section we explore what the possible function of this head is, and its connection to the 9.9 mover head.
Figure 20: Examples of high attention in examples in OpenWebText-10K for head 8.3 in GPT-2
Figure 21: Examples of high attention in examples in OpenWebText-10K for head 6.1 in GPT-2
Figure 22: Examples of high attention in examples in OpenWebText-10K for head 6.7 in GPT-2
Figure 23: We look at the top scoring component matrices in every head before layer 9 to attention head 9.9. We find several heads not previously identified in circuit analysis that have strong communication to it. These are attention heads 6.1.1, 6.2.5, 6.7.3, and 8.3.0. 8.3.0 in particular, has the strongest composition with 9.9 out of any head.
Figure 24: Per component composition scores for the 4 identified heads communicating with 9.9. Each has an outlier component that communicates most strongly with 9.9’s QK circuit.
I.2.2 Attention Head 8.3 as a Relevant Context Head
The attention patterns in Figure 20 suggest this head plays some role in attending to relevant continuations of the current context. The strong composition to the 9.9 mover head QK circuit (Figure 23) also suggests it is modulating what the copying head attends to. But what is the actual function of this head? First, we establish that this head is not a mover or inhibition head, then provide evidence that it is a head involved in signaling relevant context that can be copied to the next token by the mover head (9.9). We begin to outline how exactly the 8.3.0 component matrix interacts with the attention pattern of 9.9 with some success, but can not make strong conclusions because we can not control for every variable in the data we have available. We hope this work encourages future study on this head’s interactions with mover heads as well as future work in static weight analysis more generally.
8.3 is not a mover head
It’s possible that 8.3 acts like a mover head, copying the tokens it attends to into the residual stream, which incidentally causes 9.9 to do the same. We perform the IOI copying head test from Wang et al. [2022] and find that this is likely not the case. In this test, the output of an attention head is decoded into the vocab space. If the top token from this decoding is the same as the token it attends to, then it is “copying" that token. We find that 8.3 gets 0% on this test, while 9.9 gets 100%. Other inhibition heads also get 0%, so next we explore whether 8.3 could be interpreted as a type of inhibition head.
8.3 is not an inhibition head
We include 8.3.0 (the zeroth component in 8.3, which strongly composes with 9.9) in our analysis in Figure 11, and find it does not act like an inhibition head on the IOI task (although, as we could expect from the composition score, it does modulate the inhibition scores). Additionally, removing the zeroth component reduces the inhibition score on IOI by less than .01 (same experiment as performed in Figure 2). From these experiments, we can not rule out that 8.3 never acts like an inhibition, i.e., it could do so on a distribution of text we do not test, but evidence seems to point against this. Regardless, we can provide evidence for this head being involved with identifying relevant continuations (rather than preventing repetitions) in the next sections
8.3 attends to relevant continuations of the current context
Based on the observational evidence from the attention patterns on excerpts from OpenWebText-10K, we hypothesize that 8.3 will selectively attend to tokens that are logical continuations of the current text. To test this, we create a small synthetic counterfactual dataset of verbs and nouns to test 8.3’s attentional preferences. Our intuition is that: given some mention of a noun $n$ in context (“pork"), and a verb $v$ later in context (“eat"), 8.3 will only attend to $n$ if it is an appropriate direct object of $v$. Therefore, in the contexts “I see the pork. I am eating" and “…I am drinking", 8.3 will attend from eating to “pork" strongly, but not from “drinking" to “pork", since this is not a typical object of the verb “drink". We generate 84 counterfactual pairs of a verb and a likely/unlikely object using the verbs and nouns in Table 1. An unlikely object is any noun that does not appear in the verb’s list of likely objects. We find that 8.3 attends much more strongly to nouns that are likely to follow the given verb. Results are shown in Figure 25.
Verbs Likely Objects
cook [carrot, cabbage, steak, pork, chicken]
eat [carrot, cabbage, steak, pork, chicken]
drink [water, juice, milk, soda, coffee]
read [book, newspaper, magazine, comic, blog]
watch [movie, TV, video, show, cartoon]
open [door, window, box, jar, can]
write [letter, note, script, report]
play [game, piano, guitar, violin, drum, song]
paint [wall, fence, door, floor, ceiling]
Table 1: Verbs and appropriate objects used to test 8.3’s attention patterns to relevant objects
Figure 25: When predicting the continuation of sentences like ‘I saw the book. I am reading the’, head 8.3 is attends much more strongly to the given noun (book) when the verb is appropriate vs. inappropriate (e.g., eating). The average attention score is around 35% for likely nouns and around 5% for unlikely nouns.
Causal Interventions on 8.3.0
Next, we use the output space of the 8.3.0 component as a steering vector (since it outputs onto a line) to change the attention of 9.9, as predicted by the composition score to 9.9’s QK circuit. We use excerpts from documents in OpenWebText-10K that highly activate 8.3 (Figure 20). We hypothesize that 8.3 plays some role in telling 9.9 to attend and copy to relevant context through its zeroth component matrix. Specifically, we test whether adding some constant times the 8.3.0 output vector to the residual stream will cause 9.9 to attend more or less to the token 8.3 attends to (following a similar methodology as in Figure 4). For instance, given the input “would be better to negotiate with David, to give [him] some money and to [allow]", 8.3 attends from “allow" to “him". We find that this does tend to be the case, although it is difficult to quantitatively measure. Our results in Figure 26 show per-example attention scores after adding 8.3.0 multiplied by -1000 to the residual stream. We see a slight increase on average, but the results are inconsistent across examples. One reason is because of multitoken phrases; e.g., for phrases like “the corridor", 8.3 will attend consistently to the last token in the phrase, here “corridor". But if 9.9 is to copy this phrase, it should attend to “the", not corridor. Our test does not cover this, and it isn’t clear how to fairly handle each edge case. We suggest that some interaction like the one we’ve described here is taking place, but leave more rigorous analysis on the exact interactions between 8.3 and 9.9 for further study.
Figure 26: Interventions on 8.3.0 cause an increase in the attention of 9.9 onto the token 8.3 attends, although inconsistently. The x-axis shows results for individual examples. This is an average increase from 9% to 12%.
Appendix J More Information on Composition
J.1 Value Composition: Duplicate Token Heads
@@ I.2 Discovering New Connections
J.2 Query Composition: Inhibition Heads
@@
J.3 Key Composition: Induction Heads
@@
Appendix K Extra Laundry List Interventions
The results for scaling individual components across Laundry List datasets (varying number of objects) are in Figures 28, 29, 30, 31, 32, 33, 34, 35, and 36